Yesterday morning (March 9th), we performed some measurements to investigate the origin of the 5Hz comb which contaminates the B1p quadrant demodulated signals on SDB2.
1/First measurements with ITF in LN3:
- We switched off the cameras power, according to the following sequence:
2020-03-09-07h31m38-UTC info masserot Power camera SDB2_B1p_Cam off from mezzanine SDB2_MezzCAM
2020-03-09-07h33m44-UTC info masserot Power camera SDB2_B5_Cam off from mezzanine SDB2_MezzCAM
2020-03-09-07h34m50-UTC info masserot Power camera SDB2_B1_Cam off from mezzanine SDB2_MezzCAM
2020-03-09-07h35m52-UTC info masserot Power camera SDB2_B1s2_Cam off from mezzanine SDB2_MezzCAM
2020-03-09-07h37m30-UTC info masserot Power camera SDB2_ElectronicBox off from mezzanine SDB2_MezzCAM
- We stopped the imaging server DET_Img_CEB at 07h39m48 and took data for 4 minutes.
- We commented all CAMERA_TRIGGER in SDB2_dbox_bench and reconfigured the camera mezzanine at 07h49m08. We then collected data for 30 minutes.
- The power supply of the SDB2 switch was switched off at 08h24m51, and switched back on at 08h42m13. Soon after the ITF unlocked.
2/ Measurements with ITF unlocked:
- At about 08h55 utc, Nicolas stopped the LED on the DAQ boxes of SDB2. Starting from ~09h05 utc, the following mezzanines have also their LED switched off: DAQ box 7, mez service 2 ; DAQ box 9 mez 1 (ADC) ; DAQ box 11, mez0 et mez1 (services).
- Starting from 09h20m21, all SDB2 cameras were enabled again.
- DET_Img_CEB server started at 09h20m36 and stopped at 09h25m45. During this period, around 09h23 Nicolas desacticated the LED on DBOX 26 LNFS.
- DET_Img_CEB server started at 09h29m28 and stopped at 09h32m34.
- All cameras were commented in DET_Img_CEB and server started at 09h38m38. Taking data up to 09h41m33.
- B1p camera enabled in DET_Img_CEB and server restarted at 09h41m41. Taking data up to 09h43m32.
- B1 camera enabled in DET_Img_CEB and server restarted at 09h43m41. Taking data up to 09h46m01.
- B5 camera enabled in DET_Img_CEB and server restarted at 09h46m10. Taking data up to 09h48m09.
- B1s2 camera enabled in DET_Img_CEB and server restarted at 09h48m20.
- The power supply of the SDB2 switch was switched off at 09h51m06, and switched back on at 09h55m13.
3/ Measurements with ITF in LN3 and SDB2 LEDs switched off:
- All cameras commented in DAQ box configuration file and camera mezzanine reconfigured at 10h30m38.
- SDB2_ElectronicBox camera restored and camera mezzanine reconfigured at 10h37m01.
- B1 camera restored and camera mezzanine reconfigured at 10h41m36.
- B5 camera restored and camera mezzanine reconfigured at 10h47m31.
- B1p camera restored and camera mezzanine reconfigured at 10h52m06.
- B1s2 camera restored and camera mezzanine reconfigured at 10h57m38.
- All cameras restored at 11h02m07.
Analysis of these data is in progress. The results will be posted later.
At the end of the shift, the LEDs of SDB2 electronic boards are left off.